For each of the following situations involving single amounts, solve for the unknown. Assume that interest is compounded annually. (i = interest rate, and n = number of years) (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided. Round your final answers to nearest whole dollar amount.)

Present Value Future Value i n
1. $78,000 12.0% 6
2. $29,002 $92,000 15
3. $19,084 $49,500 10.0%
4. $51,349 $195,000 14
5. $28,012 12.0% 10
Answer #1
Answer Table Used and Workings
1 $ 39,517 PV of $1:-$78,000*PVIF(12%,6) =$78,000*0.50663 =$39,517
2 8% FV of $1:-$29,0002*FVIF(8%,15) =$29,002*3.1722 =$92,000
3 10 Years FV of $1:-$19,084*FVIF(10%,10) =$19,084*2.5937 =$49,500
4 10% FV of $1:-$51,349*FVIF(10%,14) =$51,349*3.7975 =$195,000
5 $ 87,000 FV of $1:-$28,012*FVIF(12%,10) =$51,349*3.1058 =$87,000